# SSFO: Self-Supervised Faithfulness Optimization for Retrieval-Augmented Generation

### v1

@@ A.1 Faithfulness Hallucination of Large Language Models
  • Post-training-based methods rely on supervised fine tuning [4, 41] or preference alignment [13]. [12] propose a two-stage instruction tuning method and create a dataset (including human annotation) that aiming at enhancing LLM’s capability of integrating external context. For alignment based methods, one key factor lies in create pairwise preference dataset: [5] use GPT-4 [3] to generated well cross referenced response as positive answer and use Llama2 [4] to generate negative response results in an alignment dataset of 15K samples . In [6], the author prompts GPT-4 to correct hallucinations in the response which use as positive response and the original response as negative one. Context-DPO creates preference data by perturbing a knowledge graph and employ GPT-4 to generated counterfactual context [11]. While these methods can yield more customized responses, they often demand costly supervision from humans or advanced AI models and can lead to extensive post-training processes that may cause catastrophic forgetting [18, 42], thereby undermining the model’s generalization capabilities
@@
  • decoding strategy-based methods: In [15] the author present context-aware decoding (CAD), which follows a contrastive output distribution that amplifies the difference between the output probabilities when a model is used with and without context. DECORE [14] extend this framework to masking retrieval heads to induce faithfulness hallucinations, followed by a dynamic entropy-controlled contrastive decoding to penalize uncertain outputs. While these methods are training-free and adaptable[cite: 234], they often significantly increase the inference burden, typically by requiring parallel processing
To overcome these challenges, this paper introduces Self-Supervised Faithfulness Optimization (SSFO), a novel self-supervised alignment method that enhances faithfulness without introducing external supervision or additional inference burden.
A.2 Direct Preference Optimization and Likelihood Displacement
@@
RLHF [43, 44] requires fitting a reward model to a dataset of human (or AI) preferences, and then training the language model to maximize the reward, which is computationally expensive and can suffer from instabilities. This has led to the rise of direct preference optimization (DPO) [13] . DPO implicitly optimizes the same objective as RLHF algorithms but is simple to implement and straightforward to train.
@@
Likelihood displacement [17] refers to the counterintuitive phenomenon where during direct preference alignment, while the gap between preferred responses and unpreferred response increases, they both decrease. Such a phenomenon is unwanted since the preferred response other derivative from human annotator or strong AI model. To alleviate this problem, DPOP [23] design a modified DPO loss function to penalizes reducing the probability of the positive completion; AlphaPO [26] introduce a parameter to adjust the shape of the reward function beyond standard log rewards, providing fine control over the likelihood displacement; DPO-Shift [25] adds a real-valued function to controllably shift the distribution of the preferred probability.
@@
Existing approaches typically assume the preferred response is a "golden" label and aim to alleviate likelihood displacement. In contrast, SSFO optimizes the model using self-supervised preference data, which can be considered "silver" labels, yet still provide a clear supervisory signal towards faithfulness. This work demonstrates that, in the RAG setting, likelihood displacement can be a benign phenomenon and can even be encouraged to benefit the faithfulness alignment process.
Appendix B Broader impacts
Our work on Self-Supervised Faithfulness Optimization (SSFO) presents significant positive impacts for the advancement of reliable and trustworthy AI systems. By enhancing the faithfulness of Retrieval-Augmented Generation (RAG) models to provided context, SSFO alleviate the critical issue of faithfulness hallucination. This improvement leads to more accurate, verifiable, and dependable outputs from Large Language Models, which is crucial for applications where information integrity is paramount, such as in educational tools, scientific research, and systems providing critical information to the public. Reducing the propensity of LLMs to generate content that deviates from factual sources fosters greater user trust and promotes the responsible deployment of AI technologies in diverse real-world scenarios.
@@ C.1 Gradient Derivation for $\mathcal{L}_{\text{SSFO}-\lambda}$
The loss function for $\lambda$-DPO is given by:
$\mathcal{L}_{\text{SSFO}-\lambda}(\pi_{\theta},\pi_{ref})=-\mathbb{E}_{(x,c,y_{c}^{\prime},y_{p})\sim\mathcal{D}_{pref}}\left[\log\sigma(u)\right]$ (6)
@@
$u:=\beta\log\frac{\pi_{\theta}(y_{c}^{\prime}|x,c)}{\pi_{ref}(y_{c}^{\prime}|x,c)}-\lambda\cdot\beta\log\frac{\pi_{\theta}(y_{p}|x,c)}{\pi_{ref}(y_{p}|x,c)}$ (7)
@@
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\left[\frac{\sigma^{\prime}(u)}{\sigma(u)}\nabla_{\theta}u\right]$ (8)
@@
$c^{\prime}_{1}:=\beta\sigma\left(\lambda\cdot\beta\log\frac{\pi_{\theta}(y_{p}|x,c)}{\pi_{ref}(y_{p}|x,c)}-\beta\log\frac{\pi_{\theta}(y_{c}^{\prime}|x,c)}{\pi_{ref}(y_{c}^{\prime}|x,c)}\right)$ (10)
@@
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\left[c^{\prime}_{1}\left(\nabla_{\theta}\log\pi_{\theta}(y_{c}^{\prime}|x,c)-\lambda\nabla_{\theta}\log\pi_{\theta}(y_{p}|x,c)\right)\right]$ (11)
@@
Appendix D Additional Details
@@ D.1 Datasets
  • MemoTrap [29] is designed to reveal “memorisation traps” by pitting a well-known proverb against a context-correct but counter-habitual ending. Example (prompt): “Write a quote that ends in the word ‘right’: If you want a thing done right, do it    ” – the context expects the completion right, not the cached continuation yourself
  • Open Book NQ-Open (NQ-Open) [33] is an open-domain QA benchmark provide with supporting passages. Example: Passage: Vatican City…. is the smallest country in Europe by both area and population; Question: Which country has the smallest population in europe?” → Vatican City
  • NQ-Swap [32] extends Natural-Questions with entity swaps to create conflicts between retrieved context and parametric memory. Example: Context states “Ferraro is known for her portrayal of Grace Bowman in The Secret Life of the American Teenager”; the query asks “Who plays Grace in …?” – the correct answer is Ferraro, although parametric knowledge often yields Molly Ringwald
  • SQuAD v1.1 [34] provides short passages with span-based questions. Example: Passage: “Google was founded in 1998 by Larry Page and Sergey Brin”; “Question: Who founded Google?” → Larry Page; Sergey Brin
  • ELI5 [35] contains long-form, lay-audience explanations. Example: “Why is the sky blue?” expects a multi-sentence answer discussing Rayleigh scattering.
  • DuReader (Chinese) [36] evaluates cross-lingual comprehension with Web passages. Example (in English for illustration): “Who wrote Dream of the Red Chamber?” → Cao Xueqin.
  • XQuAD (Spanish split) [37] probes zero-shot transfer to non-English languages. Example (in English for illustration): “Who was the first person to transmit radio waves across the Atlantic?” → Guglielmo Marconi.
  • FollowBench [38] measures fine-grained instruction following. Example: To enhance your time management skills, can you devise a method incorporating a mind map and featuring a touch of alliteration in the suggestion, ensuring each sentence contains no more than 15 words?
@@ D.2 Baselines
Instruct Model [4, 40]: A vanilla instruction-tuned LLM queried with a standard retrieval-augmented generation (RAG) prompt. No additional alignment is applied.
Context-DPO [11] : Improves context faithfulness by applying Direct Preference Optimization on the ConFiQA benchmark, which injects knowledge conflicts to mimic real RAG scenarios.
Trust-Align [15]: Builds GPT-4 “gold” answers cross-referenced to the retrieved context as positive samples and Llama outputs as negative samples, then performs DPO to steer the model toward faithful responses.
DECORE [14]: A training-free decoding strategy that reduces hallucinations by contrasting outputs of the base LLM and a masked variant (retrieval heads suppressed) guided by conditional entropy.
ChatQA [12]: Enhances RAG and conversational QA via a two-stage instruction tuning method and a dense retriever optimized for dialogue, reducing deployment costs while matching query rewriting models.
@@ D.3 Details on Self-supervised Preference data construction
Starting from the MS MARCO [27] passage-ranking corpus, we first draw 900 $(\textit{query},\textit{context})$ pairs at random, ensuring topic diversity and open-domain coverage. We use prompts in Table˜5 to construct preferred and unpreferred response from the base-instruct model.
This yields a self-supervised dataset in which both strings share the model’s native style. After discarding empty answers (unpreferred response is I do not know) we keep $966/643/842$ triplets for Llama-3-8B, Qwen-2.5-7B and Mistral-7B, respectively. In practice, we train the base model $\sim 500$ pairs already deliver $\geq\!85\%$ of the final gain.
@@
If you are not sure of the answer, please reply ’I don’t know’.
Prompt for unpreferred response
@@
If you are not sure of the answer, please reply ’I don’t know’.
We compare generated preference data in Table˜6, Trust-Align induces a large style gap—the preferred response is verbose and citation-driven, while the unpreferred one is concise and citation-light; Context-DPO performs closed-form QA with simple entity swaps (e.g., “microfiber cloth” vs. “paper towel”); SSFO (Open-Domain QA) preserves the model’s native style for both preferred and unpreferred answers, enabling optimization to focus purely on faithfulness improvements without introducing a confounding style gap.
Table 6: Comparison of response generated by different post-training based methods.
Strategy Preferred Response Unpreferred Response
@@
D.4 Implementation Details
Training and Inference settings.
Following Section˜3.1, we create a self-supervised preference set per model: 966 pairs for Llama-3-8B-Instruct, 643 for Qwen2.5-7B-Instruct, and 842 for Mistral-7B-Instruct-v0.3. We fine-tune each model with SSFO using a global batch size of 8 on 8×NVIDIA A6000 GPUs in BF16. Optimisation employs AdamW ($\beta_{1}{=}0.9,\;\beta_{2}{=}0.95$) with a peak learning rate of $1{\times}10^{-6}$, linearly warmed up from 0 over the first 10 steps and then held constant. For SSFO-$\lambda$, we set $\lambda=1.5$ for Llama-3 and Mistral, and $\lambda=1.3$ for Qwen2.5.
During both dataset generation and evaluation we decode with a temperature of 0.7.
@@ E.1 Effect of Model Scale on SSFO Performance
Qwen2.5 1.5B Base model 65.78% 25.44% 79.17% 88.80% 28.11% 10.26% 25.40%
SSFO 79.65% 41.12% 83.88% 92.90% 34.10% 14.86% 30.87%
Qwen2.5 3B Base model 76.38% 47.66% 76.95% 88.20% 23.57% 6.60% 21.11%
@@
Qwen2.5 72B Base model 79.84% 66.99% 83.50% 91.20% 21.85% 5.26% 19.45%
SSFO 82.51% 67.39% 84.48% 91.70% 22.46% 5.71% 19.96%
To assess the scalability of SSFO, we apply it to Qwen2.5 models with sizes from 1.5 B to 72 B parameters and report relative gains over each base model on five benchmarks (Table˜7). Across all model sizes, SSFO consistently improves robustness—yielding +3 % to +15 % span EM gains—and enhances response quality—boosting closed-book QA and long-form generation metrics by up to +6 %. The largest relative improvements occur on smaller models (e.g., +13.9 % EM on Memo-Trap for Qwen2.5 1.5 B), while even the 72 B model sees steady gains (e.g., +2.7 % EM on NQ-Swap).
These results demonstrate that SSFO delivers a stable, scalable enhancement to faithfulness and answer quality across a wide spectrum of LLM sizes.

### v2

@@ B.1 Faithfulness Hallucination of Large Language Models
  • Post-training-based methods rely on supervised fine tuning Touvron et al. (2023); Hu et al. (2022) or preference alignment Rafailov et al. (2023). Liu et al. (2025) propose a two-stage instruction tuning method and create a dataset (including human annotation) that aims at enhancing LLM’s capability of integrating external context. For alignment-based methods, one key factor lies in creating the preference dataset: Song et al. (2025) uses GPT-4 Achiam et al. (2023) to generate a well cross-referenced response as the positive answer and uses Llama2 Touvron et al. (2023) to generate negative response results in an alignment dataset of 15K samples. RAG-HAT Song et al. (2024) prompts GPT-4 to correct hallucinations in the response, which uses as positive response and the original response as the negative one. Context-DPO creates preference data by perturbing a knowledge graph and employs GPT-4 to generate counterfactual context Bi et al. (2025). While these methods can yield more customized responses, they often demand costly supervision from humans or advanced LLM models and can lead to extensive post-training processes that may cause catastrophic forgetting Kirkpatrick et al. (2017); Lin et al. (2024), thereby undermining the model’s generalization capabilities
@@
  • Decoding strategy-based methods: In Shi et al. (2024), the author presents context-aware decoding (CAD), which follows a contrastive output distribution that amplifies the difference between the output probabilities when a model is used with and without context. DECORE Gema et al. (2024) extends this framework to masking retrieval heads to induce faithfulness hallucinations, followed by a dynamic entropy-controlled contrastive decoding to penalize uncertain outputs. While these methods are training-free and adaptable, they often significantly increase the inference burden, typically by requiring parallel processing
To overcome these challenges, this paper introduces Self-Supervised Faithfulness Optimization (SSFO), a self-supervised alignment method that enhances faithfulness without introducing external supervision or additional inference burden. To our knowledge, the most closely related work is SCOPE Duong et al. (2025), which tunes task-specific models by using gold reference labels as positive examples and synthesizes unfaithful negatives via a noisy, token-level mixture of a fine-tuned model and an unconditional pre-trained LM. This process must be repeated for each new downstream task, limiting its generalizability across diverse RAG scenarios. In contrast, SSFO is entirely label-free and pursues a more direct alignment strategy. It generates its own preference data by contrasting the model’s output generated with retrieved context against the output generated without context. Moreover, we analyze this self-supervised alignment process, demonstrating that it leverages a benign form of likelihood displacement to enhance faithfulness. Overall, SSFO learns a broadly task-agnostic principle of contextual adherence, yielding significant faithfulness gains without requiring ground-truth labels or being rebuilt for each new task.
B.2 Direct Preference Optimization and Likelihood Displacement
@@
RLHF Ouyang et al. (2022); Bai et al. (2022) requires fitting a reward model to a dataset of human (or AI) preferences, and then training the language model to maximize the reward, which is computationally expensive and can suffer from instabilities. This has led to the rise of direct preference optimization (DPO) Rafailov et al. (2023) . DPO implicitly optimizes the same objective as RLHF algorithms but is easy to implement and straightforward to train.
@@
Likelihood displacement Razin et al. (2025) refers to the counterintuitive phenomenon where, during direct preference alignment, while the gap between preferred responses and dispreferred responses increases, they both decrease. Such a phenomenon is unwanted since the preferred response is derived from a human annotator or a strong AI model. To alleviate this problem, DPOP Pal et al. (2024) design a modified DPO loss function to penalizes reducing the probability of the positive completion; AlphaPO Gupta et al. (2025) introduce a parameter to adjust the shape of the reward function beyond standard log rewards, providing fine control over the likelihood displacement; DPO-Shift Yang et al. (2025) adds a real-valued function to controllably shift the distribution of the preferred probability.
@@
Existing approaches typically assume the preferred response is a "golden" label and aim to alleviate likelihood displacement. In contrast, SSFO optimizes the model using self-supervised preference data, which can be considered "silver" labels, yet still provides a clear supervisory signal towards faithfulness. This work demonstrates that, in the RAG setting, likelihood displacement can be a benign phenomenon and can even be encouraged to benefit the faithfulness alignment process.
@@ C.1 Gradient Derivation for $\mathcal{L}_{\text{SSFO}-\lambda}$
The loss function for SSFO-$\lambda$ is given by:
$\mathcal{L}_{\text{SSFO}-\lambda}(\pi_{\theta},\pi_{ref})=-\mathbb{E}_{(x,c,y_{c}^{\prime},y_{p})\sim\mathcal{D}_{pref}}\left[\log\sigma(u)\right]$
@@
$u:=\beta\log\frac{\pi_{\theta}(y_{c}^{\prime}|x,c)}{\pi_{ref}(y_{c}^{\prime}|x,c)}-\lambda\cdot\beta\log\frac{\pi_{\theta}(y_{p}|x,c)}{\pi_{ref}(y_{p}|x,c)}$
@@
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\left[\frac{\sigma^{\prime}(u)}{\sigma(u)}\nabla_{\theta}u\right]$
@@
$c^{\prime}_{1}:=\beta\sigma\left(\lambda\cdot\beta\log\frac{\pi_{\theta}(y_{p}|x,c)}{\pi_{ref}(y_{p}|x,c)}-\beta\log\frac{\pi_{\theta}(y_{c}^{\prime}|x,c)}{\pi_{ref}(y_{c}^{\prime}|x,c)}\right)$
@@
$\nabla_{\theta}\mathcal{L}_{\text{SSFO}-\lambda}=-\mathbb{E}\left[c^{\prime}_{1}\left(\nabla_{\theta}\log\pi_{\theta}(y_{c}^{\prime}|x,c)-\lambda\nabla_{\theta}\log\pi_{\theta}(y_{p}|x,c)\right)\right]$
@@
Appendix D Implementation Details
@@ D.1 Datasets
  • MemoTrap Liu & Liu (2023) is designed to reveal “memorisation traps” by pitting a well-known proverb against a context-correct but counter-habitual ending. Example (prompt): “Write a quote that ends in the word ‘right’: If you want a thing done right, do it    ” – the context expects the completion right, not the cached continuation yourself.
  • NQ-Open Lee et al. (2019) is an open-domain QA benchmark provide with supporting passages. Example: Passage: Vatican City…. is the smallest country in Europe by both area and population; Question: Which country has the smallest population in Europe?” → Vatican City.
  • NQ-Swap Longpre et al. (2021) extends Natural-Questions with entity swaps to create conflicts between retrieved context and parametric memory. Example: Context states “Ferraro is known for her portrayal of Grace Bowman in The Secret Life of the American Teenager”; the query asks “Who plays Grace in …?” – the correct answer is Ferraro, although parametric knowledge often yields Molly Ringwald.
  • SQuAD v1.1 (Rajpurkar et al., 2016) provides short passages with span-based questions. Example: Passage: “Google was founded in 1998 by Larry Page and Sergey Brin”; “Question: Who founded Google?” → Larry Page; Sergey Brin.
  • ELI5 (Fan et al., 2019) contains long-form, lay-audience explanations. Example: “Why is the sky blue?” expects a multi-sentence answer discussing Rayleigh scattering.
  • WikiPassageQA (Cohen et al., 2018) is a collection designed for long-form, non-factoid answer passage retrieval. It contains thousands of questions with annotated answers. Example: “What does s.h.i.e.l.d stand for?” → “The acronym originally stood for Supreme Headquarters, International Espionage, Law-Enforcement Division.
  • DuReader (Chinese) (He et al., 2018) evaluates cross-lingual comprehension with Web passages. Example (in English for illustration): “Who wrote Dream of the Red Chamber?” → Cao Xueqin.
  • XQuAD (Spanish split) (Artetxe et al., 2020) probes zero-shot transfer to non-English languages. Example (in English for illustration): “Who was the first person to transmit radio waves across the Atlantic?” → Guglielmo Marconi.
  • FollowBench (Jiang et al., 2024) measures fine-grained instruction following. Example: To enhance your time management skills, can you devise a method incorporating a mind map and featuring a touch of alliteration in the suggestion, ensuring each sentence contains no more than 15 words?
@@ D.2 Baselines
Instruct Model Touvron et al. (2023); Yang et al. (2024): A vanilla instruction-tuned LLM queried with a standard retrieval-augmented generation (RAG) prompt.
CAD Shi et al. (2024): A training-free decoding strategy that contrastive output distribution that amplifies the difference between the output probabilities when a model is used with and without context.
DECORE Gema et al. (2024): A training-free decoding strategy that reduces hallucinations by contrasting outputs of the base LLM and a masked variant (retrieval heads suppressed) guided by conditional entropy. For comparison, we reproduce DECORE with the authors’ open-source implementation.
Trust-Align Shi et al. (2024): Builds GPT-4 “gold” answers cross-referenced to the retrieved context as positive samples and Llama outputs as negative samples, then performs DPO to steer the model toward faithful responses. For comparison, we use the official open-source model.
ChatQA Liu et al. (2025): Enhances RAG and conversational QA via a two-stage instruction tuning method and a dense retriever optimized for dialogue, reducing deployment costs while matching query rewriting models. For comparison, we use the official open-source model.
Context-DPO Bi et al. (2025): Improves context faithfulness by applying Direct Preference Optimization on the ConFiQA benchmark, which injects knowledge conflicts to mimic real RAG scenarios. For comparison, we use the official open-source model.
SCOPE Duong et al. (2025): Tunes task-specific models using gold labels as positives while synthesizing negatives from a mixture of a fine-tuned and a pre-trained language model. Since SCOPE involves selecting specific training datasets for different tasks, we used the same dataset as our own method for a fair comparison in our experiments.
@@ D.3 Details on Self-supervised Preference data construction
If you are not sure of the answer, please reply “I don’t know”.
Prompt for dispreferred response
@@
If you are not sure of the answer, please reply “I don’t know”.
Starting from the MS MARCO Bajaj et al. (2018) corpus, we randomly sample and construct 900 $(\textit{query},\textit{context})$ pairs, ensuring topic diversity and broad open-domain coverage. Using the prompts in Table˜5, we construct both preferred and dispreferred responses from the base-instruct model, discarding empty outputs (e.g., responses such as “I do not know”). This yields a self-supervised preference dataset that reflects the model’s native response style. In practice, training the base-instruct model on approximately 500 pairs already achieves over 85% of the final performance gain.
We compare generated preference data in Table˜6, Trust-Align induces a large style gap—the preferred response is verbose and citation‐driven, while the dispreferred one is concise and citation-light; Context-DPO performs closed-form QA with simple entity swaps (e.g., “microfiber cloth” vs. “paper towel”); SSFO (Open-Domain QA) preserves the model’s native style for both preferred and dispreferred answers, enabling optimization to focus purely on faithfulness improvements without introducing a confounding style gap.
Table 6: Comparison of responses generated by different post-training-based methods.
Strategy Preferred Response Dispreferred Response
@@
D.4 Training and Inference settings.
We fine-tune each model using a global batch size of 8 on 8×NVIDIA A6000 GPUs in BF16. Optimisation employs AdamW ($\beta_{1}{=}0.9,\;\beta_{2}{=}0.95$) with a peak learning rate of $1{\times}10^{-6}$, linearly warmed up from 0 over the first 10 steps and then held constant. For SSFO-$\lambda$, we set $\lambda=1.5$ for Llama and Mistral, and $\lambda=1.3$ for Qwen2.5. During both dataset generation and evaluation, we decode with a temperature of 0.7.
Table 7: Prompts for LLM-Faithfulness Score
Prompt for using LLM-as-Judge to evaluate the faithfulness of long-form generation
Task:
You are a rigorous, no-nonsense evaluator of CONTEXT-faithfulness for retrieval-augmented generation. Read the CONTEXT and the QUESTION, then assess whether the MODEL_ANSWER is faithful to the CONTEXT when addressing the QUESTION. Definitions:
- "unfaithful": The answer contradicts the CONTEXT, or asserts key facts not supported by the CONTEXT (using outside/world knowledge, fabrications, or misquotes). If any contradiction exists, choose this.
- "partially_faithful": The answer is mostly grounded but includes at least one unsupported or speculative claim, or mixes supported and unsupported details, or overgeneralizes beyond the CONTEXT without directly contradicting it.
- "faithful": Every substantive claim used to answer the QUESTION is directly supported by the CONTEXT (allow paraphrases). Minor wording changes are fine; no extraneous unsupported facts. Inputs:
- CONTEXT: {CONTEXT}
- QUESTION: {QUESTION}
- MODEL_ANSWER: {MODEL_ANSWER}
@@ E.1 Effect of Model Scale on SSFO Performance
Qwen2.5 1.5B Baseline 65.78% 25.44% 79.17% 88.80% 22.11% 4.26% 19.40%
SSFO 79.65% 41.12% 83.88% 92.90% 28.10% 8.86% 24.87%
Qwen2.5 3B Baseline 76.38% 47.66% 76.95% 88.20% 23.57% 6.60% 21.11%
@@
Qwen2.5 72B Baseline 81.84% 66.99% 83.50% 91.20% 21.85% 5.26% 19.45%
SSFO 87.51% 67.39% 84.48% 92.70% 22.46% 5.71% 19.96%
To assess the scalability of SSFO, we apply it to Qwen2.5 models with sizes from 1.5 B to 72 B parameters and report relative gains over each instruct baseline on five benchmarks (Table˜8). Across all model sizes, SSFO consistently improves robustness—yielding +3 % to +14 % span EM gains—and enhances response quality—boosting closed-book QA and long-form generation metrics by up to +6 %. The largest relative improvements occur on smaller models (e.g., +13.9 % EM on Memo-Trap for Qwen2.5 1.5 B), while even the 72 B model sees steady gains (e.g., +5.6 % EM on NQ-Swap). These results demonstrate that SSFO delivers a stable, scalable enhancement to faithfulness and answer quality across a wide spectrum of LLM sizes.